OpenStudy (anonymous):

Solve :) What is the solution of the equation 6x − 3 = -51?

OpenStudy (anonymous):

add 3 to both sides

OpenStudy (anonymous):

A. -9 B. -8 C. 8 D. 9

OpenStudy (anonymous):

you will get 6x = -48 then divide both sides by 6 x = -8

OpenStudy (anonymous):

6x − 3 = -51 +3 +3 6x = -48 /6 /6 x = -8

OpenStudy (anonymous):

Well 3 added to each sides is what you would want to start with. Therefore, you are left with 48 = 6x which is -8 (the solution)